# TPV: Parameter Perturbations Through the Lens of Test Prediction Variance

### v1

@@ 5 Experiments
We run experiments around TPV stability between training and test sets and empirically identify the conditions when it holds and when it breaks. We then study TPV under label noise and how width plays an important role in modulating it. We do not include experiments for SGD based perturbations and the relation between the flatness of minima and TPV because several prior research works have confirmed a similar relationship (keskar2017large; smith2018bayesian; jastrzkebski2017three). Finally, we run experiments on the proposed JBR pruning algorithm comparing its performance to existing post-training pruning methods.
@@ 5.2 Empirical Evaluation of TPV Under Label Noise
Next, we increase the label-noise standard deviation to $\sigma=0.1$. This larger noise requires models to move farther from the clean reference solution, violating the first-order approximation underlying Eq. 10. As expected, TPV stability breaks: the training-set empirical TPV approaches $\sigma^{2}$ once the noisy model fits the training data, and the dependence on geometric structure diminishes. However, the test-set empirical TPV remains aligned with the theoretical TPV (particularly at larger widths), and continues to correlate with test loss (Figure 7). Thus, even when the first-order assumption is violated, test-set TPV continues to reflect generalization behavior.
@@ 6 Discussion and conclusion
We introduced TPV as a simple but powerful variance-based framework for probing the robustness of a trained neural network under specific noise sources. We proved that in overparameterized regime, training-set TPV tends to test-set TPV, and empirically showed that this TPV stability holds remarkably well even in extremely narrow networks, as long as a sufficient number of samples is used and the first order approximation holds. This leads us to believe that we may have found a sufficient condition under which TPV stability holds, but it may not be a necessary one.
@@
TPV can be used to measure the robustness of a trained model w.r.t. specific perturbation sources and therefore acts as a surgical tool to study generalization at a fine-grained level. In this paper, we showed that studying TPV for both SGD noise (Theorem 3.3) and quantization noise (Remark 3.4) leads to the wide minima hypothesis. This hypothesis is widely used as a tool for model selection and even guiding optimization (foret2020sharpness). Under the TPV framework, we see that this hypothesis probes the robustness of a model under a specific class of parameter perturbations. Modeling label noise as a source of perturbation under the TPV framework leads to a different TPV estimate (Theorem 3.2) that depends on the spectrum of the output-parameter Jacobian and its alignment between training and test data distributions, and revealed a different practical approach to measuring TPV. We empirically found this object to also correlate with the standard test set loss for various datasets and models. One can similarly probe TPV under perturbations induced by a different noise source.
@@ C.2 Useful Regime and Pathological Case Analysis
In this section we resolve two subtle but important conceptual points regarding: i) the regime in which the label-noise TPV (Theorem 3.2) is useful; ii) pathological cases where the label-noise TPV object are seemingly different for the training and test sets and TPV stability appears to breaks.
@@
If now, the test-set TPV from Theorem 3.2 still depends on the Jacobian spectrum and therefore varies across architectures, it may raise the concern that TPV stability has failed. We show below that this is not the case: the interpolation scenario is a special, degenerate limit that violates the assumptions of Theorem 3.2, and is in fact incompatible. If on the other hand, the training and test distributions are identical, then test-set TPV becomes identical to the training set TPV ($\sigma_{\varepsilon}^{2}$). In this case, TPV stability holds trivially, but the label noise TPV object itself loses its discriminative power and is no longer informative.
To shed light on the above concerns, we begin by noting that Theorem 3.2 does not assume interpolation of noisy labels. Theorem 3.2 analyzes the local linearized problem around the clean minimizer $w^{\star}$. Let $J\in\mathbb{R}^{n\times p}$ be the Jacobian of network outputs at the training inputs. Under label noise $\varepsilon$, the linearized loss is
@@
(3) Low-rank or effectively narrow networks: If $r<n$, then even with $\lambda=0$, the interpolation residual $\varepsilon_{\mathrm{res}}=(I-U_{r}U_{r}^{\top})\varepsilon$ is non-zero, and
@@
(4) Large-noise, non-local regime (Figure 7): There remains one additional regime, which we empirically probe in Fig. 7: the large-noise, non-local regime. For sufficiently large noise variance $\sigma_{\varepsilon}^{2}$, SGD driven by the noisy logits may move the parameters far away from $w^{\star}$. In this case: (i) the parameter increment $\delta w$ is no longer small, (ii) the Jacobian $J(w)$ along the optimization path changes substantially, and (iii) the actual solution reached by SGD is not well-approximated by the local min-(ridge-)norm solution of the linearized problem at $w^{\star}$.
Empirically, in this regime we observe that the training noise is effectively interpolated (the network fits the large noisy perturbations), so the empirical training TPV saturates near $\sigma_{\varepsilon}^{2}$ and becomes almost independent of width, while the test TPV still varies with width. At first glance this seems to violate TPV stability. However, this behavior occurs precisely when the assumptions of Theorem 3.2 are violated: the first-order approximation around $w^{\star}$ breaks down and the min-(ridge-)norm linearized dynamics no longer describe the optimization. In other words, the large-noise experiment in Fig. 7 is intentionally outside the small-perturbation regime where our theory applies. As a result, the mismatch between empirical train and test TPV in this experiment does not constitute a counterexample to TPV stability; it merely marks the breakdown of the underlying linearization and optimization assumptions.
@@
On the other hand, the concern that “expressive models always interpolate the noisy logits, forcing $\operatorname{TPV}_{\text{train}}=\sigma_{\varepsilon}^{2}$ while test TPV varies” arises only when implicitly combining: (i) full row rank ($r=n$) due to exact interpolation on training set, (ii) zero regularization ($\lambda=0$), and (iii) ignoring the breakdown of the first-order approximation at large noise. Theorem 3.2 assumes min-norm dynamics within the local linear region around $w^{\star}$. In the regime where these assumptions hold (small perturbations, local solutions, matched train/test distributions), both training and test TPV depend on the Jacobian geometry and TPV stability makes training-set TPV a useful object. Note that if the first order approximation is not violated but the other two conditions are still true, then, if training and test set distributions are identical, test-set TPV has to be equal to the training-set TPV because of the mathematical relation shown in the case 2a above. In the extreme large-noise or shifted-distribution regimes (e.g., Fig. 7), the theory is not expected to apply, and empirical violations of TPV stability simply signal the breakdown of these assumptions rather than a flaw in the underlying results.
The practically relevant regime, and the one probed in our experiments, is the nontrivial, small-but-finite perturbation setting where the optimization remains local, the noisy labels are not interpolated in the linear approximation. In this regime, TPV retains meaningful dependence on the Jacobian geometry as predicted by Theorem 3.2.
@@ C.3.4 Practical Resolution: Local SGD With Proximity Regularization
    Models need to be trained in eval mode: we perturb the logits of the clean model’s prediction with Gaussian noise and train a copy of the reference model to fit these new targets, which act as infinitesimal change in targets. To achieve this faithfully, training must be done in eval mode, i.e., modules like batch norm and dropout should not be present. The reason is that if these modules are in training mode, even a zero variance perturbation causes the model to have perturbed targets, which is not our goal.
@@
    Mini-batch shuffling: All sources of randomness other than label noise should be removed as much as possible to isolate the effect of label noise when measuring TPV (e.g. different mini-batch in each epoch). In practice, we do you mini-batch SGD for noisy label fine-tuning for efficiency and to make the training loss go down in some cases. However, we ensure that the sequence of mini-batch remains the same in each epoch or at the very least across different runs. Even though mini-batch gradient is a sum of the expected gradient plus a noise vector, removing random shuffling in this way ensures that each independent run sees the same sequence of mini-batches.
@@ D.1 Relation Between Gradient Covariance Matrix and Hessian near Minima
Note that similar proportionality between SGD gradient covariance and curvature (Hessian / Fisher) has been shown in prior work (wu2022alignment; mori2022powerlaw; mandt2017variational; kuehn2023correlated; barshan2020fisherhessian), typically under population or online assumptions, specific model classes (linear networks, RFMs) or log-likelihood objectives. Here we make this connection explicit for finite-sample nonlinear squared-loss regression on a fixed dataset with no label noise.
@@ F.3 Pruning Experiment Details
This appendix provides the experimental details for the pruning results reported in Section 5.3. All pruning experiments follow the global structured pruning protocol introduced in Optimal Brain Compression (OBC) (chen2025optimal). Our implementation is based on the official GitHub code of (chen2025optimal) and extends the OBC framework by adding the TPV-motivated JBR importance criteria.
@@ F.3.1 Pruning Framework
We adopt the global channel-pruning pipeline of chen2025optimal. At each pruning iteration, an importance score is computed for every pruning group (typically convolutional or linear output channels, together with the corresponding input-channel dependencies). The least-important groups are removed globally, and the dependency graph ensures architectural consistency. No fine-tuning is performed between pruning iterations.
@@ F.3.3 Pruning Criteria Compared
  • Jacobian (chen2025optimal),
  • L1 weight-norm (Li2017Pruning),
@@
  • BatchNorm scale (liu2017learning),
  • FPGM (he2019filter) ,
  • WHC (chen2023whc),
  • Taylor first-order saliency (molchanov2019importance).

### v2

@@ 5 Experiments
We run experiments around TPV stability between training and test sets and empirically identify the conditions when it holds and when it breaks. We then study TPV under label noise and how width plays an important role in modulating it. We do not include experiments for SGD based perturbations and the relation between the flatness of minima and TPV because several prior research works have confirmed a similar relationship (Keskar and et al., 2017; Smith and Le, 2018; Jastrzebski et al., 2017). Finally, we run experiments on the proposed JBR pruning algorithm comparing its performance to existing post-training pruning methods.
@@ 5.2 Empirical Evaluation of TPV Under Label Noise
Next, we increase the label-noise standard deviation to $\sigma=0.1$. In this case, we find that TPV stability breaks and training and test set TPV behave differently– training-set empirical TPV approaches $\sigma^{2}$ once the noisy model fits the training data, and the dependence on geometric structure diminishes, while the test-set empirical TPV remains aligned with the theoretical TPV (particularly at larger widths), and continues to correlate with test loss (Figure 7). We conjecture that this happens because $\sigma=0.1$ induces a large perturbation covariance which weakens the upper bound in Theorem 2.1, diminishing the TPV stability guarantee.
@@ 6 Discussion and conclusion
We introduced TPV as a simple but powerful variance-based framework for probing the robustness of a trained neural network under specific noise sources. We proved that in overparameterized regime, training-set TPV tends to test-set TPV, irrespective of the model’s generalization performance, and empirically showed that this TPV stability holds remarkably well even in extremely narrow networks, as long as a sufficient number of samples is used, and the induced parameter perturbation covariance is small. This leads us to believe that Theorem 2.1 may be a sufficient condition under which TPV stability holds, but it may not be a necessary one.
@@
TPV can be used to measure the robustness of a trained model w.r.t. specific perturbation sources and therefore acts as a surgical tool to study generalization at a fine-grained level. In this paper, we showed that studying TPV for both SGD noise (Theorem 3.3) and quantization noise (Remark 3.4) leads to the wide minima hypothesis. This hypothesis is widely used as a tool for model selection and even guiding optimization (Foret et al., 2020). Under the TPV framework, we see that this hypothesis probes the robustness of a model under a specific class of parameter perturbations. Modeling label noise as a source of perturbation under the TPV framework leads to a different TPV estimate (Theorem 3.2) that depends on the spectrum of the output-parameter Jacobian and its alignment between training and test data distributions, and revealed a different practical approach to measuring TPV. We empirically found this object to also correlate with the standard test set loss for various datasets and models. One can similarly probe TPV under perturbations induced by a different noise source.
@@ C.2 Useful Regime and Pathological Case Analysis
In this section we resolve two subtle but important conceptual points regarding: i) the regime in which TPV stability holds and the label-noise TPV (Theorem 3.2) is useful; ii) pathological cases where label noise perturbations are too large and and TPV stability breaks.
@@
If now, the test-set TPV from Theorem 3.2 still depends on the Jacobian spectrum and therefore varies across architectures, it may raise the concern that TPV stability has failed. We argue below that this case happens when the label noise is too large to the point that either the first order assumption breaks (in which case neither Theorem 2.1 nor Theorem 3.2 apply), or the upper bound of the distance between training and test set TPV in Theorem 2.1 becomes too loose (since it directly depends on the trace of the perturbation covariance) and TPV stability does not hold based on the theorem.
If on the other hand, the label noise is not too large but the training and test distributions are identical, then test-set TPV becomes identical to the training set TPV ($\sigma_{\varepsilon}^{2}$). In this case, TPV stability holds trivially, but the label noise TPV object itself loses its discriminative power and is no longer informative.
To shed light on the above cases, we begin by noting that Theorem 3.2 does not assume interpolation of noisy labels. Theorem 3.2 analyzes the local linearized problem around the clean minimizer $w^{\star}$. Let $J\in\mathbb{R}^{n\times p}$ be the Jacobian of network outputs at the training inputs. Under label noise $\varepsilon$, the linearized loss is
@@
(3) Low-rank or effectively narrow networks, and small perturbations: If $r<n$, then even with $\lambda=0$, the interpolation residual $\varepsilon_{\mathrm{res}}=(I-U_{r}U_{r}^{\top})\varepsilon$ is non-zero, and
@@
(4) Large-noise (Figure 7): There remains one additional regime, which we empirically probe in Fig. 7: the large-noise regime. For sufficiently large noise variance $\sigma_{\varepsilon}^{2}$, SGD driven by the noisy logits may move the parameters far away from $w^{\star}$. In this case, the parameter increment $\delta w$ is no longer small. This can lead to at least one of two situations: (i) the trace of the induced parameter covariance matrix is large, (ii) the actual solution reached by SGD is not well-approximated by the local min-(ridge-)norm solution of the linearized problem at $w^{\star}$ and the first order approximation breaks.
Empirically, in Figure 7, we observe that the training noise is effectively interpolated (the network fits the large noisy perturbations), so the empirical training TPV saturates near $\sigma_{\varepsilon}^{2}$ and becomes almost independent of width, while the test TPV still varies with width, but closely matches the theoretical TPV. Thus, while TPV stability breaks, theorem 3.2 holds. Since theorem 3.2 requires the first order approximation to hold, we conjecture that TPV stability breaks because the large noise induces parameter perturbations whose covariance matrix trace is large, which in turn weakens the upper bound in the TPV trace stability theorem 2.1 and diminishes the TPV stability guarantee.
@@
On the other hand, if the parameter perturbations induced by label noise is too large, it can make the upper bound in the TPV stability theorem weak, or worse, break the first order assumption. This diminishes TPV stability guarantee. In this situation, if the first order assumption holds, theorem 3.2 remains applicable (Figure 7), however, TPV stability nonetheless fails and training set cannot be used for estimating robustness.
The practically relevant regime, and the one probed in our experiments (e.g. Figure 5), is the nontrivial, small perturbation setting where the optimization remains local, TPV stability holds, and the noisy labels are not interpolated in the linear approximation. In this regime, TPV retains meaningful dependence on the Jacobian geometry as predicted by Theorem 3.2 and training set estimate of TPV remains a good estimator of the test set estimate.
@@ C.3.4 Practical Resolution: Local SGD With Proximity Regularization
    Models need to be trained in eval mode: we perturb the logits of the clean model’s prediction with Gaussian noise and train a copy of the reference model to fit these new targets, which act as infinitesimal change in targets. To achieve this faithfully, training must be done in eval mode, i.e., modules like batch norm and dropout should not be active. The reason is that if these modules are in training mode, even a zero variance perturbation causes the model to have perturbed targets, which conflicts with our goal.
@@
    Mini-batch shuffling: All sources of randomness other than label noise should be removed as much as possible to isolate the effect of label noise when measuring TPV (e.g. different mini-batch in each epoch). In practice, we do use mini-batch SGD for noisy label fine-tuning for efficiency and to make the training loss go down in some cases. However, we ensure that the sequence of mini-batch remains the same in each epoch or at the very least across different runs. Even though mini-batch gradient is a sum of the expected gradient plus a noise vector, removing random shuffling in this way ensures that each independent run sees the same sequence of mini-batches.
@@ D.1 Relation Between Gradient Covariance Matrix and Hessian near Minima
Note that similar proportionality between SGD gradient covariance and curvature (Hessian / Fisher) has been shown in prior work (Wu et al., 2022; Mori, 2022; Mandt et al., 2016; Kühn and Rosenow, 2023; Barshan et al., 2020), typically under population or online assumptions, specific model classes (linear networks, RFMs) or log-likelihood objectives. Here we make this connection explicit for finite-sample nonlinear squared-loss regression on a fixed dataset with no label noise.
@@ F.3 Pruning Experiment Details
This appendix provides the experimental details for the pruning results reported in Section 5.3. All pruning experiments follow the global structured pruning protocol introduced in Optimal Brain Compression (OBC) (Chen et al., 2025). Our implementation is based on the official GitHub code of (Chen et al., 2025) and extends the OBC framework by adding the TPV-motivated JBR importance criteria.
@@ F.3.1 Pruning Framework
We adopt the global channel-pruning pipeline of Chen et al. (2025). At each pruning iteration, an importance score is computed for every pruning group (typically convolutional or linear output channels, together with the corresponding input-channel dependencies). The least-important groups are removed globally, and the dependency graph ensures architectural consistency. No fine-tuning is performed between pruning iterations.
@@ F.3.3 Pruning Criteria Compared
  • Jacobian (Chen et al., 2025),
  • L1 weight-norm (Li et al., 2017),
@@
  • BatchNorm scale (Liu et al., 2017),
  • FPGM (He et al., 2019) ,
  • WHC (Chen et al., 2023),
  • Taylor first-order saliency (Molchanov et al., 2019).